The sum of the lengths of one side of a right triangle and the hypotenuse
is 20 cm. The other leg is 5 cm shorter than the hypotenuse. Find the
length of each side.

Answer by scott8148(5892)

let x=hypotenuse, so 20-x="one side" and x-5="other side"

by Pythagoras, (20-x)^2+(x-5)^2=x^2 ___ 400-40x+x^2+x^2-10x+25=x^2

x^2-50x+425=0 ___ use quadratic formula to find x